# The Baryonic Tully-Fisher Relation of HI-bearing Low Surface Brightness Galaxies Implies Their Formation Mechanism

### v2

@@ 4 Results
where $\alpha$ and $\beta$ are both free parameters, to fit the BTFRs of the HSBGs and LSBGs, respectively. The best-fitting parameters of the BTFR of HSBGs are $\alpha_{\rm HSBG}=1.83\pm 0.51$ and $\beta_{\rm HSBG}=3.91\pm 0.24$, while the best-fitting parameters of LSBGs are $\alpha_{\rm LSBG}=-1.45\pm 1.49$ and $\beta_{\rm LSBG}=5.67\pm 0.78$.
We further split LSBGs and HSBGs into the different $\log v_{\rm{rot}}$ bins. Within each bin, we calculate the median value of $\log M_{\rm{b}}$ and the corresponding $1\sigma$ standard deviation for both samples. As shown in Figure 1, the binned points of LSBGs (depicted by the blue squares), as well as the binned points of HSBGs (depicted as the black squares), comfortably lie within the $1\sigma$ uncertainty region of the tightest BTFR of typical late-type galaxies reported by Lelli et al. (2019) (the red-solid line and the red-shaded region denote the best-fitting BTFR of typical late-type galaxies and its $1\sigma$ uncertainty, respectively, as shown in Fig. 1), with $\beta\simeq 3.85\pm 0.09$.
@@
Figure 1: $v_{\rm rot}-M_{\rm b}$ diagram for our samples. The green circles represent LSBGs, while the grey circles represent HSBGs, and the orange circles are SPARC galaxies from Lelli et al. (2019). All the error bars denote $\rm 1\sigma$ uncertainties. The blue squares represent the median values of LSBGs while the black squares represent that of HSBGs. The red-solid line is the tightest BTFR reported by Lelli et al. (2019). The red-shaded area is the $\rm 1\sigma$ intrinsic scatter of the red-solid line, which is considered as the $\rm 1\sigma$ uncertainty of the tightest BTFR. The sub-panels display the number-density distribution of LSBGs (green histograms) and HSBGs (grey histograms).
@@ 5 Discussion
5.1 Selection biases
In our study, the SNR threshold used to select the LSBG and HSBG samples may exclude some galaxies with similar HI masses but wider HI line widths from our parent sample, which might potentially introduce deviations in the slope of BTFR. To investigate the possible selection effects, we adjust the SNR threshold by increasing it from 10 to 15, or 20.
We find that, for HSBGs, the best-fitting $\beta$ of their BTFR changes to $4.05\pm 0.21$ for SNR $>15$, and $4.05\pm 0.26$ for SNR $>20$, still in good agreement with the best-fitting $\beta$ value for SNR$>10$. For LSBGs, the best-fitting $\beta$ becomes $\beta\sim 6.04\pm 0.66$ (SNR $>15$) and $\sim 6.13\pm 0.82$ (SNR $>20$), which is also similar to the $\beta\sim 5.67\pm 0.78$ when applying the selection criterion of SNR$>10$, considering the large uncertainties of the $\beta$ values.
Additionally, the binned points of LSBGs and HSBGs within the different $\log v_{\rm{rot}}$ for the different SNR thresholds are always located within the $1\sigma$ uncertainty region of the BTFR of typical late-type galaxies. Consequently, we conclude that, for both of LSBG and HSBG samples, the selection effect introduced by the SNR threshold is negligible.
5.2 The radial acceleration relation
Apart from BTFR, recently, a series of studies (McGaugh et al., 2016; Lelli et al., 2017; Li et al., 2018; Richards et al., 2018; Rong et al., 2018; Chae et al., 2020) have revealed another tight correlation known as the radial acceleration relation (RAR). RAR indicates that the acceleration $g_{\rm bar}$ produced by the baryonic mass and the acceleration $g_{\rm obs}$ produced by the dynamical mass at any radius in many types of galaxies consistently follow a correlation of
$g_{\rm obs}=\frac{g_{\rm bar}}{1-e^{-\sqrt{g_{\rm bar}/a_{0}}}},$ (12)
where $a_{0}\simeq(1.20\pm 0.02\pm 0.24)\times 10^{-10}\ {\rm m/s^{2}}$ (McGaugh et al., 2016; Lelli et al., 2017) is constant. In the $\rm\Lambda CDM$ paradigm, RAR plausibly suggests a universally established link between the mass distributions of baryonic matter and dark matter in galaxies.
We estimate $g_{\rm obs}$ and $g_{\rm bar}$ of each galaxy at its HI-radius ($R_{\rm HI}$) using the following equations,
$g_{\rm obs}=\frac{v_{\rm rot}^{2}}{R_{\rm HI}},$ (13)
$g_{\rm bar}=x_{\rm*}\frac{GM_{\rm*}}{R^{2}_{\rm HI}}+x_{\rm g}\frac{GM_{\rm g}}{R^{2}_{\rm HI}},$ (14)
where $G$ is the gravitational constant. $x_{\rm*}$ and $x_{\rm g}$ in equation (14) are correction factors assuming that the cold gas and stars are distributed in two razor-thin disks, respectively. They are calculated as (Freeman, 1970),
$x_{\rm*}={\rm I_{0}}(\frac{R_{\rm HI}}{2R_{d,*}}){\rm K_{0}}(\frac{R_{\rm HI}}{2R_{d,*}})-{\rm I_{1}}(\frac{R_{\rm HI}}{2R_{d,*}}){\rm K_{1}}(\frac{R_{\rm HI}}{2R_{d,*}}),$ (15)
$x_{\rm g}={\rm I_{0}}(\frac{R_{\rm HI}}{2R_{d,g}}){\rm K_{0}}(\frac{R_{\rm HI}}{2R_{d,g}})-{\rm I_{1}}(\frac{R_{\rm HI}}{2R_{d,g}}){\rm K_{1}}(\frac{R_{\rm HI}}{2R_{d,g}}),$ (16)
where $\rm I_{n}$ and $\rm K_{n}$ are the modified Bessel functions of the first and second kind, respectively; $R_{\rm d,*}$ and $R_{\rm g,*}$ involved are the scale lengths of the stellar and HI disks which can be derived by using the same method described in the work of Rong et al. (2024).
$R_{\rm HI}$ corresponds to the outer region of a galaxy where the surface density of HI equals $1\ M_{\rm\odot}/{\rm pc^{2}}$ (Broeils & Rhee, 1997). $R_{\rm HI}$ can be estimated from the following tight correlation between HI mass and $R_{\rm HI}$, which has been observed in various types of galaxies including LSBGs (Wang et al., 2016; Gault et al., 2021; Lutz et al., 2018),
$\log(\frac{2R_{\rm HI}}{\rm kpc})=(0.506\pm 0.003)\log(\frac{M_{\rm HI}}{M_{\rm\odot}})-(3.293\pm 0.009).$ (17)
In Figure 2, at $R_{\rm HI}$, both our sample LSBGs and HSBGs follow the universal RAR (McGaugh et al., 2016; Lelli et al., 2017, as illustrated by the red components in Fig. 2). The similarity of RARs between LSBG and HSBGs may suggest comparable mass distributions in the two galaxy samples.
It is worth noting that, at a large radius of a galaxy where $g_{\rm bar}\ll a_{0}$, such as $R_{\rm{HI}}$, the RAR described by equation (12) is actually equivalent to the BTFR with $\beta=4$ (Lelli et al., 2017).
Figure 2: The RAR of our sample galaxies, overplotted by data from Lelli et al. (2016a). The circles and squares are the same as those in Figure 1. The red-solid line is the best-fitted RAR in previous studies (McGaugh et al., 2016; Lelli et al., 2017), and the grey-dot-dashed line is the identical line.
@@ 5.3 Implications on the formation mechanisms of LSBGs
Previous theoretical studies posit that LSBGs may form in low-density dark matter halos (e.g. Dekel & Silk, 1986; McGaugh, 1992). However, this model predicts a significant difference in the BTFRs between LSBGs and HSBGs (McGaugh, 2021), which contradicts our findings. Therefore, the formation model based on low-density dark matter halos may not refer to LSBGs.
Many observational studies also suggest that LSBGs might be dominated by dark matter compared with HSBGs (e.g. de Blok & McGaugh, 1997). Therefore, we delve deeper into understanding the fraction of dark matter in LSBGs. Note that BTFR could be expressed as  [^1] Equation (19) is from the lecture on Tully-Fisher relation; see also Aaronson et al. (1979) and Zwaan et al. (1995).,
@@
where $R$ corresponds to a large radius in the outer region of a galaxy (such as $R_{\rm{HI}}$), and
$f_{\rm b}(R)=\frac{M_{\rm b}}{M_{\rm b}+M_{\rm DM}(R)},$ (19)
is the baryonic mass fraction within $R$, where $M_{\rm{DM}}(R)$ denotes the dark matter mass enclosed within $R$. $g_{\rm bar}(R)$ is the baryonic acceleration at $R$, which could be calculated with equation (14). Consequently, $f_{\rm b}(R_{\rm{HI}})$ of the LSBGs and HSBGs can be estimated with equation (18).
@@
We examine the variation of $f_{\rm b}(R_{\rm HI})$ as a function of $v_{\rm{rot}}$ for both LSBGs and HSBGs. As shown in Fig. 3, in statistics, $f_{\rm b}(R_{\rm HI})$ increases with rising $v_{\rm{rot}}$, with a slope of $\sim 1.02$ using a bisector fitting (Isobe et al., 1990). Since the dark matter halo mass of a galaxy is tightly related to $v_{\rm{rot}}$, as $M_{\rm{h}}\propto v_{\rm{rot}}^{3}$ (Mo et al., 1998), the correlation between $f_{\rm b}(R_{\rm HI})$ and $v_{\rm{rot}}$ hints at a positive association between the baryonic mass fractions and dark matter halo masses for these galaxies (Ayromlou et al., 2023; Duffy et al., 2010). This correlation does not exceed our expectations, given that our galaxy samples are dominated by the low-mass galaxies. Halos with lower $M_{\rm{h}}$ have shallower potentials (Duffy et al., 2010), making it more difficult to prevent the galactic interstellar medium (ISM) against expulsion by the feedback processes in these shallower potentials, which leads to lower $f_{\rm b}(R_{\rm{HI}})$ (e.g. Bryan et al., 2013; Duffy et al., 2010).
@@
Subsequently, in order to compare the $f_{\rm b}(R_{\rm{HI}})$ of the LSBG and HSBG samples with similar dark matter halo masses, we further focus on a specific $v_{\rm{rot}}$ bin defined by $\log v_{\rm rot}\in[1.87,2.05]$ (as depicted by the blue-shaded region in Fig. 3). We have 52 LSBGs and 52 HSBGs in this small bin, exhibiting similar distributions of $v_{\rm rot}$ with median $\log v_{\rm rot}$ values of $1.94\pm 0.05$ (for LSBGs) and $1.97\pm 0.05$ (for HSBGs). The Kuiper test between the distributions of $v_{\rm rot}$ of LSBGs and HSBGs in this small bin also gives a large $p$-value of $0.77$, indicating no significant difference in $v_{\rm rot}$. We then compare the $f_{\rm b}(R_{\rm HI})$ distributions of the LSBG and HSBG sample galaxies within this $v_{\rm rot}$ bin. The median values of the $f_{\rm b}(R_{\rm HI})$ distributions of LSBGs and HSBGs are $-0.94\pm 0.15$ and $-0.91\pm 0.30$, respectively; the Kuiper test for the two $f_{\rm b}(R_{\rm HI})$ distributions also yields a large $p$-value of $0.34$.
@@
Therefore, we note that, for the LSBGs and HSBGs with similar halo masses, their baryonic fractions are also comparable. LSBGs are not particularly dark matter dominated galaxies, compared with the HSBG counterparts. Furthermore, this result also plausibly suggests comparable feedback power among the LSBGs and HSBGs with similar halo masses (i.e., similar gravitational potentials). Consequently, we conclude that feedback may not be the primary driver of the different surface brightness between these HI-bearing galaxies. In other words, these HI-bearing LSBGs may not originate from stronger or weaker feedback.
@@
Alternatively, our finding aligns with the simulation results where the LSBGs can be well reproduced by higher dark matter halo spins (Pérez-Montaño et al., 2022). Therefore, we conclude that our result may lend support to the theorem that LSBGs may originate from high-spin dark matter halos (Mo et al., 1998).
Figure 3: The $v_{\rm rot}-f_{\rm b}(R_{\rm HI})$ relation of our sample galaxies. The circles are the same as those in Figure 1. The blue-shaded area represents the bin in which we perform Kuiper tests in Section 5.3. The black-dashed line is the best-fitted line for the data points.
5.4 Implications on MOND
Except for the aforementioned formation models of LSBGs in the framework of $\rm\Lambda CDM$, MOND may also offer a reasonable explanation for the BTFR and RAR of LSBGs (e.g. Milgrom, 1983; Wittenburg et al., 2020).
As demonstrated in Section 4 and 5.2, both LSBGs and HSBGs adhere to a BTFR with a slope indistinguishable from 4, and may follow the same RAR as given by equation (12). These results adhere to the theoretical predictions by MOND (Milgrom, 1983). Therefore, our result cannot exclude the MOND theory.

### v3

@@ 4 BTFR
where $\alpha$ and $\beta$ are free parameters to fit the BTFRs of the HSBGs and LSBGs using scipy.odr package, respectively. To estimate the errors of $\alpha$ and $\beta$, we apply the Monte Carlo method. We generate data samples based on the original ($v_{\rm rot}$, $M_{\rm b}$) values and their associated uncertainties. We assume Gaussian distributions for $v_{\rm rot}$ and $M_{\rm b}$ centered at their origins, with the $1\sigma$ regions matching the error bars. This process yields a set of ($\alpha$, $\beta$) values. We repeat this sampling procedure 5,000 times to obtain 5,000 sets of ($\alpha$, $\beta$) values. The standard deviations of these parameters ($\sigma_{x,{\rm MC}}$) are then combined with the fitting errors ($\sigma_{x,{\rm odr}}$) to determine the uncertainties of $\alpha$ and $\beta$, as
$\sigma_{x}=\sqrt{\sigma_{x,{\rm MC}}^{2}+\sigma_{x,{\rm odr}}^{2}},$ (9)
where the subscript $x$ denotes $\alpha$ or $\beta$.
The best-fitting parameters of the BTFR are $\alpha_{\rm HSBG}=1.83\pm 0.51$ and $\beta_{\rm HSBG}=3.91\pm 0.24$ for HSBGs, while are $\alpha_{\rm LSBG}=-1.45\pm 1.49$ and $\beta_{\rm LSBG}=5.67\pm 0.78$ for LSBGs.
We further split LSBGs and HSBGs into the different $\log v_{\rm{rot}}$ bins. Within each bin, we calculate the median value of $\log M_{\rm{b}}$ and the corresponding $1\sigma$ standard deviation for both samples. As shown in Fig. 3, the binned points of LSBGs (depicted by the blue squares) and HSBGs (depicted as the black squares) comfortably lie within the $1\sigma$ uncertainty region of the tightest BTFR of typical late-type galaxies reported by Lelli et al. (2019) (the red-solid line and the red-shaded region denote the best-fitting BTFR of typical late-type galaxies and its $1\sigma$ uncertainty, respectively, as shown in Fig. 3), with $\beta\simeq 3.85\pm 0.09$.
@@
Our results are consistent with many previous observational (e.g., McGaugh, 2012; Lelli et al., 2016a; Papastergis et al., 2016) and simulation (e.g., Dutton et al., 2017) results. However, it is also worth noting that some other studies alternatively report smaller slopes ($\beta\sim 3$, e.g., Goddy et al., 2023; Ponomareva et al., 2018).
Figure 3: The $v_{\rm rot}-M_{\rm b}$ diagram for our samples. The green- and the grey-filled circles are sthe same as those in Figure 2. The orange-filled circles are SPARC galaxies from Lelli et al. (2019). All the error bars denote $\rm 1\sigma$ uncertainties. The blue squares represent the median values of LSBGs while the black squares represent those of HSBGs. The red-solid line is the tightest BTFR reported by Lelli et al. (2019). The red-shaded area is the $\rm 1\sigma$ intrinsic scatter of the red-solid line, which is considered as the $\rm 1\sigma$ uncertainty of the tightest BTFR.
@@ 5 Discussion
5.1 Potential biases
In our study, the SNR threshold used to select the LSBG and HSBG samples may exclude some galaxies with similar HI masses but wider HI line widths, potentially introducing deviations in the slope of the BTFR. To investigate the possible selection effects, we adjust the SNR threshold by increasing it from 10 to 15, and then to 20. We find that, for HSBGs, the best-fitting $\beta$ of their BTFR changes to $4.05\pm 0.21$ for SNR $>15$, and $4.05\pm 0.26$ for SNR $>20$, still in good agreement with the best-fitting $\beta$ value for SNR $>10$. For LSBGs, the best-fitting $\beta$ becomes $\sim 6.04\pm 0.66$ (SNR $>15$) and $\sim 6.13\pm 0.82$ (SNR $>20$), which is also similar to the $\beta\sim 5.67\pm 0.78$ when applying the selection criterion of SNR $>10$, considering the large uncertainties of the $\beta$ values. Additionally, the binned points of LSBGs and HSBGs within different $\log v_{\rm{rot}}$ bins for the different SNR thresholds are always located within the $1\sigma$ uncertainty region of the BTFR of typical late-type galaxies. Consequently, we conclude that, for both the LSBG and HSBG samples, the selection effect introduced by the SNR threshold is negligible. It should also be noted that, due to the sensitivity limitations of the Arecibo telescope, galaxies with higher rotational velocities $v_{\rm rot}$ exhibit lower signal-to-noise ratios in their HI spectra under a fixed total HI flux, leading to a bias in our parent sample (Haynes et al., 2011; Dou et al., 2024).
Another potential bias may arise from the $k_{4}$ cut. Some rotation-supported galaxies with small $v_{\rm rot}$ may exhibit single-horned spectra due to beam-smearing effects (Swaters et al., 2003; Di Teodoro & Fraternali, 2015), leading to their exclusion by the $k_{4}$ cut. Additionally, galaxies with larger inclination angles tend to have smaller $k_{4}$ values (El-Badry et al., 2018). We study the effect of the $k_{4}$ cut by adjusting the $k_{4}$ threshold from -1.0 to -1.2, and then from -1.0 to -0.8. The BTFR slope of LSBGs changes from $\sim 5.67\pm 0.78$ ($k_{4}<-1.0$) to $\sim 5.38\pm 0.70$ ($k_{4}<-1.2$) and $\sim 4.28\pm 0.32$ ($k_{4}<-0.8$), while the slope of HSBGs changes from $\sim 3.91\pm 0.24$ ($k_{4}<-1.0$) to $\sim 3.78\pm 0.24$ ($k_{4}<-1.2$) and $\sim 3.85\pm 0.16$ ($k_{4}<-0.8$), respectively. It is worth noting that when the $k_{4}$ threshold exceeds $-0.8$, some galaxies appear to deviate from the BTFR. However, these galaxies may be strongly dominated by velocity dispersions, as they typically have unreliable $v_{\rm rot}\lesssim 40\ {\rm km/s}$ (Lelli, 2022; El-Badry et al., 2018). Since we cannot determine whether the single-horned HI spectra of these galaxies are due to the beam-smearing effect or the dominance of the velocity dispersions, we keep $k_{4}<-1$ to exclude these galaxies.
Additionally, possible bias emerges from using $W_{20}$ or $W_{50}$ (the $50\%$ peak width of an HI spectrum) to estimate the rotation velocities. However, we find that the average difference between the $\log v_{\rm rot}$ estimated from $W_{20}$ and $W_{50}$ is quite small, even smaller than the uncertainties of $\log v_{\rm rot}$. Therefore, using $W_{20}$ or $W_{50}$ would not significantly affect our results. Besides, we also test whether using the different methods of estimating $M_{\rm*}$ (Taylor et al., 2011) or considering the intrinsic thickness $q_{0}$ dependence on $M_{\rm*}$ (Lelli et al., 2016b) would lead to the changes of BTFR slopes of LSBGs. We observe no strong changes in BTFR slopes in these tests. Furthermore, we notice that the surface brightness limit determined by the optical survey may lead to large uncertainties in $b/a$ estimation for LSBGs (Mancera Piña et al., 2024). However, the large uncertainties should result in a larger scatter, rather than a systematic bias or a change of BTFR slope.
@@ 5.2 Implications on the formation mechanisms of LSBGs
Previous theoretical studies posit that LSBGs may form in low-density dark matter halos (e.g., Dekel & Silk, 1986; McGaugh, 1992). However, this model predicts a significant difference in the BTFRs between LSBGs and HSBGs (McGaugh, 2021), which contradicts our findings. Therefore, the formation model based on low-density dark matter halos may not refer to LSBGs.
Many observational studies also suggest that LSBGs might be dominated by dark matter compared with HSBGs (e.g., de Blok & McGaugh, 1997; Swaters et al., 2003; Mancera Piña et al., 2022; Read et al., 2017). Therefore, we delve deeper into understanding the fraction of dark matter in LSBGs. Note that BTFR could be expressed as  [^2] Equation (10) is from the lecture on the Tully-Fisher relation; see also Aaronson et al. (1979) and Zwaan et al. (1995).,
@@
Here, the baryonic mass fraction within a radius $R$ is defined as $f_{\rm b}(R)=M_{\rm b}/M_{\rm tot}(R)$, where $M_{\rm{tot}}(R)$ denotes the total mass enclosed within $R$. $g_{\rm bar}(R)$ is the baryonic acceleration at $R$. $g_{\rm bar}(R)$ is calculated by assuming the stellar and gas components in a galaxy follow the razor-thin disk distribution, using the method outlined in Read et al. (2017) (see equation (5) of Read et al. (2017) for details).
Note that equation (10) holds true only when $R$ is large. Here, we utilize the HI radius, $R_{\rm HI}$, defined as the radius where the HI surface density equals $1\ M_{\rm\odot}/{\rm pc^{2}}$. $R_{\rm HI}$ can be estimated based on the strong correlation between $R_{\rm HI}$ and $M_{\rm HI}$ (Wang et al., 2016; Gault et al., 2021; Lutz et al., 2018),
$\log\left(\frac{2R_{\rm HI}}{\rm kpc}\right)=(0.506\pm 0.003)\log\left(\frac{M_{\rm HI}}{M_{\rm\odot}}\right)-(3.293\pm 0.009).$ (11)
Consequently, we calculate $f_{\rm b}(R_{\rm HI})$ from equation (10) and then analyze the $f_{\rm b}(R_{\rm HI})\--v_{\rm rot}$ relationship for the LSBGs and HSBGs.
@@
As shown in panel (d) of Figure 2, in statistics, $f_{\rm b}(R_{\rm HI})$ increases with rising $v_{\rm{rot}}$, with a slope of $\sim 1.02$ using a bisector fitting (Isobe et al., 1990). Since the dark matter halo mass ($M_{\rm h}$) of a galaxy is tightly related to $v_{\rm{rot}}$, as $M_{\rm{h}}\propto v_{\rm{rot}}^{3}$ (Mo et al., 1998), the correlation between $f_{\rm b}(R_{\rm HI})$ and $v_{\rm{rot}}$ hints at a positive association between the baryonic mass fractions and dark matter halo masses for these galaxies (Ayromlou et al., 2023; Duffy et al., 2010). Similar trends have also been observed based on spatially resolved data (e.g., Lelli et al., 2017; Posti et al., 2019; Mancera Piña et al., 2022; Di Teodoro et al., 2023). This correlation does not exceed our expectations, given that our galaxy samples are dominated by the low-mass galaxies. Halos with lower $M_{\rm{h}}$ have shallower potentials (Duffy et al., 2010), making it more difficult to prevent the galactic interstellar medium (ISM) against expulsion by the feedback processes in these shallower potentials, which leads to lower $f_{\rm b}(R_{\rm{HI}})$ (e.g., Bryan et al., 2013; Duffy et al., 2010).
@@
Subsequently, in order to compare the $f_{\rm b}(R_{\rm{HI}})$ of the LSBG and HSBG samples with similar dark matter halo masses, we further focus on a specific $v_{\rm{rot}}$ bin defined by $\log v_{\rm rot}\in[1.87,2.05]$ (as depicted by the blue-shaded region in panel (d) of Fig. 2). We have 52 LSBGs and 52 HSBGs in this small bin, exhibiting similar distributions of $v_{\rm rot}$ with median $\log v_{\rm rot}$ values of $1.94\pm 0.05$ (for LSBGs) and $1.97\pm 0.05$ (for HSBGs). The Kuiper test between the distributions of $v_{\rm rot}$ of LSBGs and HSBGs in this small bin also gives a large $p$-value of $0.77$, indicating no significant difference in $v_{\rm rot}$. We then compare the $f_{\rm b}(R_{\rm HI})$ distributions of the LSBG and HSBG sample galaxies within this $v_{\rm rot}$ bin. The median values of the $f_{\rm b}(R_{\rm HI})$ distributions of LSBGs and HSBGs are $-0.94\pm 0.15$ and $-0.91\pm 0.30$, respectively; the Kuiper test for the two $f_{\rm b}(R_{\rm HI})$ distributions also yields a large $p$-value of $0.34$.
@@
Therefore, we note that, for LSBGs and HSBGs with similar halo masses, their baryonic fractions are also comparable. LSBGs may not be particularly dark matter-dominated galaxies compared to their HSBG counterparts; at least within the regimes of $R_{\rm HI}$, the dark matter fractions of LSBGs and HSBGs are similar. Furthermore, this result also plausibly suggests comparable feedback power among the LSBGs and HSBGs with similar halo masses (i.e., similar gravitational potentials). Consequently, we conclude that feedback may not be the primary driver of the different surface brightness between these HI-bearing galaxies. In other words, these HI-bearing LSBGs may not originate from stronger or weaker feedback.
@@
Alternatively, our finding aligns with the simulation results where the LSBGs can be well reproduced by higher dark matter halo spins (e.g., Pérez-Montaño et al., 2022; Di Cintio et al., 2019). Therefore, we conclude that our result may lend support to the theorem that LSBGs may originate from high-spin dark matter halos (e.g., Mo et al., 1998).
Except for the aforementioned formation models of LSBGs in the framework of $\rm\Lambda CDM$, MOND may also offer a reasonable explanation for the BTFR of LSBGs (e.g., Milgrom, 1983; Wittenburg et al., 2020). As demonstrated in Section 4, both LSBGs and HSBGs follow a BTFR with a slope indistinguishable from 4. This result adheres to the theoretical predictions by MOND (Milgrom, 1983). Therefore, our result cannot exclude the MOND theory. Nevertheless, several studies on the dynamics of galaxies have produced results that deviate from MOND’s predictions (e.g., Ren et al., 2019; Mercado et al., 2024; Mancera Piña et al., 2024; Khelashvili et al., 2024). Therefore, more compelling studies are needed to better determine if MOND is a promising framework.